# Humanoid Whole-Body Badminton via Multi-Stage Reinforcement Learning

### v1

@@ VII-C Supplementary Explanation for Model-based Hitting Target Generation
Figure 8: Trajectory generation. Shuttlecock trajectories are filtered to ensure interception points within the region $x\in[-0.8,0.8]m$ , $y\in[-1,0.2]m$ and $z\in[1.5,1.6]m$ for robot training.
@@
We generated 2 million trajectories, from which 19,685 met the criteria and were selected for robot training. The majority of these trajectories reached the hitting zone within a time interval of [0.8, 1.4] seconds, as shown in Fig. 10.
@@
VII-D Repeatability Accuracy of Swing
During the real-world Virtual-Target Swinging test, we assessed repeatability precision by commanding a single fixed interception target and executing 20 independent swings. The racket center was tracked in real time using a motion-capture system. At the designated hit instant, positional error was defined as the distance between the captured racket-center coordinate and the target position. The resulting 3D trajectories of the racket center across the 20 trials are presented in Fig. 11, while Fig. 12 visualizes the distance between the racket center and the target position over time.

### v2

@@ VII-C Supplementary Explanation for Model-based Hitting Target Generation
Figure 8: Trajectory generation. Shuttlecock trajectories are filtered to ensure interception points within the region $x\in[-0.8,0.8]\,\text{m}$, $y\in[-1,0.2]\,\text{m}$ and $z\in[1.5,1.6]\,\text{m}$ for robot training.
@@
We generated 2 million trajectories, from which 196,940 met the criteria and were selected for robot training. The majority of these trajectories reached the hitting zone within a time interval of [0.8, 1.4] seconds, as shown in Fig. 10.
@@
VII-D Sensitivity Analysis of Prediction Parameters
For the 20 real-world badminton trajectories collected in our experiments, we perform a simple sensitivity analysis with respect to the aerodynamic characteristic length $L$. For each trajectory, we randomly sample a multiplicative perturbation factor $\alpha\sim\mathcal{U}(0.8,1.2)$ and set the effective characteristic length to $L^{\prime}=\alpha L$. We then recompute the shuttle predictions under these perturbed parameters, and summarize the resulting errors in Fig. 11. Although the perturbation degrades prediction accuracy, the mean position error still converges to within 10 cm by about 0.45 s before interception.
Figure 11: EKF prediction accuracy under varying aerodynamic characteristic lengths. The parameter $\alpha$ scales the characteristic length to emulate variations in shuttle aerodynamics.
VII-E Repeatability Accuracy of Swing
During the real-world Virtual-Target Swinging test, we assessed repeatability precision by commanding a single fixed interception target and executing 20 independent swings. The racket center was tracked in real time using a motion-capture system. At the designated hit instant, positional error was defined as the distance between the captured racket-center coordinate and the target position. The resulting 3D trajectories of the racket center across the 20 trials are presented in Fig. 12, while Fig. 13 visualizes the distance between the racket center and the target position over 20 repeated swings.
@@ VII-E Repeatability Accuracy of Swing
VII-F Striking Motion Analysis
The complete real-world striking sequence for a representative return is illustrated in Fig. 14. When no shuttle is coming, the robot keeps stepping in place with the racket held in front of the body. Once a shuttle is launched, the controller initiates approach step(s) while simultaneously performing a backswing. During the hitting instant, the legs and arm accelerate in a coordinated manner to generate a whole-body impulse and produce a high racket speed. Notably, the non-racket arm swings in the opposite direction during the fast stroke, which helps counteract overall angular momentum and maintain balance. After striking, the robot performs a recovery motion.
Figure 14: Real-world striking motion. Three snapshots of a successful return in the real world. Left: in-place stepping before the shuttle is launched. Middle: approach and backswing phase as the shuttle launched (yellow box highlights the shuttle, the green arrow indicates the racket swing direction, and the white arrow indicates the stepping motion). Right: hit and follow-through: the robot simultaneously takes a step and swings the racket toward the shuttle, then completes the motion with a follow-through.